# Optimization of a cosmic muon tomography scanner for cargo border control inspection

### v1

@@ Abstract
The past several decades have seen significant advancement in applications using cosmic-ray muons for tomography scanning of unknown objects. One of the most promising developments is the application of this technique in border security for the inspection of cargo inside trucks and sea containers in order to search for hazardous and illicit hidden materials. This work focuses on the optimization studies for a muon tomography system similar to that being developed within the framework of the ‘SilentBorder’ project funded by the EU Horizon 2020 scheme. Current studies are directed toward optimizing the detector module design, following two complementary approaches. The first leverages TomOpt, a Python-based end-to-end software that employs differentiable programming to optimize scattering tomography detector configurations. While TomOpt inherently supports gradient-based optimization, a Bayesian Optimization module is introduced to better handle scenarios with noisy objective functions, particularly in image reconstruction-driven optimization tasks. The second optimization strategy relies on detailed GEANT4-based simulations, which, while more computationally intensive, offer higher physical fidelity. These simulations are also employed to study the impact of incorporating secondary particle information alongside cosmic muons for improved material discrimination. This paper presents the current status and results obtained from these optimization studies.
@@ I Introducton
Muon tomography, also known as muography, is a non-invasive imaging technique that leverages naturally occurring cosmic-ray muons to probe the internal structure of large objects. These muons originate from interactions between high-energy cosmic rays and atomic nuclei in the upper atmosphere, producing showers of secondary particles, primarily pions and kaons. As these mesons decay, they generate muons that travel toward the Earth’s surface with minimal energy loss. At sea level, the flux of cosmic-ray muons is approximately , with an average energy of vertical muons of the order . Due to their relatively large mass—roughly times that of an electron - these highly energetic muons have a strong penetrating ability, allowing them to traverse several meters of dense material. This makes them particularly valuable for imaging applications where conventional radiation-based methods, such as X-ray imaging, are ineffective due to limited penetration depth [Borozdin2003, ].
@@
Muon tomography can be classified into two main approaches: muon transmission imaging and muon scattering tomography (MST). The transmission method relies on detecting attenuation in muon flux as particles pass through an object, providing insights into its internal composition. This technique is particularly useful for large-scale applications, such as studying volcanic structures [Alessandro2018, ] or uncovering hidden chambers in archaeological sites [Alvarez1970, ]. However, it typically requires prolonged data acquisition periods, which makes it less suitable for applications where rapid imaging is essential. On the other hand, muon scattering tomography (MST) operates by measuring the deflections of muon trajectories caused by multiple Coulomb scattering interactions with the material they pass through. High-atomic-number (Z) materials, such as lead and uranium, tend to cause more significant scattering than lower-Z materials like aluminum or plastic. This property makes MST particularly useful for detecting dense and high-Z substances in diverse applications that span many fields, including nuclear reactor inspection, civil engineering, archaeology, and various industrial and homeland security uses [IAEA2022, ]. In border security applications, MST is a powerful tool for identifying hidden threats and dense contraband in cargo containers, such as radioactive materials or explosives. Furthermore, MST is well-suited for scanning smaller objects and offers the advantage of rapid data acquisition, making it an ideal technology for real-time security applications where both speed and precision are critical for border security operations [barnes2023cosmic, ].
@@
As muons traverse a material medium, they primarily interact with material nuclei through multiple Coulomb scatterings, where small-angle deviations are most frequent, while occasional large-angle deflections occur as described by Rutherford scattering. After crossing a macroscopic distance, the overall scattering angle distribution consists of a Gaussian core, which accounts for the majority of deflections due to the central limit theorem, along with a longer tail from less frequent large-angle scatterings. For a muon with momentum [MeV] passing through a material of radiation length [m] over a distance [m], the root mean square (RMS) width of the Gaussian component of the scattering angle distribution can be approximated by [pdg, ]:
@@
Optimizing the design parameters of an MST scanner is crucial to achieving optimal performance, as different configurations impact key factors such as acceptance, detection efficiency, and angular resolution. Design parameters such as the spatial arrangement and size of detectors play a critical role in determining the system’s ability to accurately track muon trajectories, maintain a high signal-to-noise ratio, and effectively classify materials. A suboptimal design may lead to reduced tracking precision, lower detection rates, and decreased overall scanning effectiveness.
In this work, we focus on optimizing the SilentBorder scanner, a modular MST system designed for cargo screening applications (FIG. 1). The scanner consists of multiple hodoscopes, each of which is a self-contained detector module comprising three equally-spaced sensitive detector panels inside a protective casing. These hodoscopes are arranged around the scanned cargo, forming a tracking system capable of measuring muon trajectories before and after interaction with the inspected volume. Due to its modular design, the SilentBorder scanner offers flexibility in configuration, allowing different setups to be explored for improved detection performance. To systematically optimize the design parameterization of the hodoscopes, we adopt a dual-method approach combining Monte Carlo simulations and differentiable optimization techniques [MODE2022, ,MODE2023, ]:
GEANT4-based simulation framework: We employ a GEANT4 [GEANT4, ] simulation framework to model different SilentBorder scanner configurations and various cargo materials. GEANT4, a widely used toolkit for simulating particle interactions with matter, provides a realistic modeling environment for evaluating the performance of different scanner layouts.
TomOpt-based optimization: TomOpt [TomOpt-Strong2023, ] is a differentiable simulation framework developed for optimizing MST scanner designs through gradient descent.
@@ II GEANT4 Simulation Framework
A GEANT4 framework was developed by the SilentBorder consortium to provide a simplified model of the SilentBorder scanner concept with the flexibility to simulate a variety of different scanner configurations and cargo contents.
Geometry Description Markup Language (GDML) [gdml, ] is used to describe the size, shape, and location of the relevant objects and materials. Firstly, an empty shipping container of external dimensions 6.07 m2.44 m2.60 m is placed at the center of the geometry with a plywood floor of depth 40 cm. The container is surrounded by a total of 48 hodoscopes, a 34 grid on each of the long sides of the container. This arrangement is visualized in FIG. 2. Each hodoscope is itself comprised of 3 layers (plates) of solid plastic scintillator each of dimensions 2.1 m1.1 m3.3 mm and separated by 10 cm. Note the casing and electronics for each hodoscope are not simulated in this simplification, but the 10 cm gaps are introduced to account for the ‘dead regions’ they introduce. It is also worth noting that the plate is simulated as solid scintillator rather than as a mat of scintillator fibers.
@@
Air shower cosmic-ray muons can be simulated in this volume using the MUons for SIlent BOrder (MuSiBo) generator. MuSiBo utilizes a Gaisser parameterization of the muon flux, with modifications to account for muon decay and the Earth’s curvature and is based on work presented in reference [Blackwell_2015, ]. Muons are sampled on the surface of a cube, a method which better accounts for the high-angle flux when compared to sampling on a single surface above the region of interest. For this simulation, a box of 11 m5 m4 m was used to fully enclose the volume of the container and all of the hodoscopes.
Finally, the initial muons simulated with MuSiBo are propagated through the simulated geometry via use of the GEANT4 software package [GEANT4, ]. Specifically, the QGSP_BERT physics list is utilized in version 11.0.p03 of the software. Muons, and any secondary particles produced, are stepped through the full geometry with information saved about their trajectory and total energy deposition within each logical volume.
@@ III GEANT4 Study of Hodoscope Positioning
The positioning of hodoscope plates with respect to one another can impact the performance of the scanner system. This impact can be appreciated in both the efficiency in collecting ‘interesting’ muons and the angular resolution of the reconstructed muons.
@@
A study is performed using the GEANT4 simulation framework described in Section II. A baseline configuration was established with the container filled entirely with wood and the hodoscopes as close together in or (in-plane) as possible. Note that due to space reserved for electronics and casing this still leaves a 10 cm gap between the active scintillator areas of adjacent hodoscopes. The default spacing in (out-of-plane) between the three plates that form a single hodoscope is 10 cm.
@@
A number of scenarios are created by independently varying the spacing in , and . Due to physical constraints (such as not overlapping any hodoscope regions) the variations are different for each direction. The cargo is kept identical and for each scenario a 15 minute exposure to the cosmic flux is provided by the MuSiBo generator.
The energy depositions recorded by GEANT4 in the active hodoscope plates pass through a pseudo detector simulation. Any depositions recorded within 1 mm and 20 ns of each other are combined and a Gaussian positional smearing, with width 1 mm, is then applied. Finally, a threshold of 0.4 MeV is applied to ensure the deposition would have created a detector ‘hit’.
@@
These hits form the input to a simple reconstruction chain. Tracks are formed when one hit is found on at least two hodoscope plates. These two plates must be found in different layers of the same hodoscope group (above, below, left or right). If a hodoscope plate has received multiple hits it is excluded from track building. Linear regression is used to find the best fit line between the hits providing the direction of the reconstructed muon track. The final stage of the reconstruction takes events in which the muon had tracks reconstructed in two different hodoscope groups. A simple Point-of-Closest-Approach (PoCA) reconstruction is performed from which the location and angle of the scatter point is determined. An example of the output of the PoCA reconstruction is shown in FIG. 3.
@@
The number of events in which a PoCA point was able to be reconstructed.
The number of events in which a PoCA point was able to be reconstructed inside the container - thus providing information about the cargo.
@@
In both directions the number of muons of interest increase as the gaps increase, this is purely because the hodoscopes as an ensemble now cover a larger volume within the total simulation world and hence more of the muons are likely to intercept at least one hodoscope. However, the number of PoCA points, and specifically those in the cargo, reduce due to the gaps that are created in the regions closest to the cargo itself. There is a subsequent reduction in the number of muons that intercept hodoscopes before and after the container and pass through the cargo. The reduction in useful PoCA points per muon will mean that a longer exposure time is required in order to achieve the same level of sensitivity.
@@
There is no visible change in the standard deviation of the angular difference metric, so the angular resolution does not improve. There is a small increase in the mean when going to the largest gaps. This is likely due to the reduced available phase space in the scattering angle created by the gaps. Moderate angles will not be measurable in these scenarios due to muons passing through the gaps between the hodoscopes.
@@
FIG. 6 shows the variation in efficiency and angular resolution metrics as a function of the vertical spacing between the hodoscope plates in . The nominal gap between the plates is 10 cm. As with the variations, the number of interesting muons increases as the total scanner volume increases with larger spacing. The number of PoCA points slightly reduces. This occurs because the probability of the muon hitting enough plates to ensure track reconstruction before and after the cargo decreases, but the effect is very minimal. The angular resolution on the other hand is dramatically impacted by the widening of the vertical spacing. By increasing the spacing both the standard deviation and mean parameters reduce significantly before beginning to flatten out. As the distance between the two or three detection points increases the positional resolution of each plate has a smaller impact on the reconstructed angle, hence improving the angular resolution. The reduction in the mean occurs due to the asymmetry of the distribution. As the resolution improves the asymmetry is naturally reduced in magnitude.
@@
These variation studies can be used to inform decisions about the design for a full-scale SilentBorder-like scanner system. As would be anticipated, minimizing the in-plane gaps between hodoscopes improves the efficiency of the system, thus reducing the necessary exposure time and making the system quicker to use and more viable in border security. Increasing the vertical out-of-plane spacing between hodoscope plates improves the angular resolution due to the lever arm effect, although it does so at the expense of efficiency. For a particular system an appropriate trade-off balance can be found, in the scenario examined in this study FIG. 6 suggests that up to 20 cm the increase in out-of-plane spacing improves the angular resolution enough for the reduction in efficiency to be worthwhile. Beyond this, the resolution improvements plateau and therefore would be unlikely to be worth the efficiency reduction. Note that these studies were completed with a simple PoCA reconstruction and decisions were made to ignore plates with multiple hits or events with 3 or more tracks. These considerations will have an impact on the magnitude of the effects presented but should not undermine the trajectories of the changes and thus the conclusions drawn.

### v2

@@ Abstract
The past several decades have seen significant advancement in applications using cosmic-ray muons for tomography scanning of unknown objects. One of the most promising developments is the application of this technique in border security for the inspection of cargo inside trucks and sea containers in order to search for hazardous and illicit hidden materials. This work focuses on the optimization studies for a muon tomography system similar to that being developed within the framework of the ‘SilentBorder’ project funded by the EU Horizon 2020 scheme. Current studies are directed toward optimizing the detector module design, following two complementary approaches. The first leverages TomOpt, a Python-based end-to-end software that employs differentiable programming to optimize scattering tomography detector configurations. While TomOpt inherently supports gradient-based optimization, a Bayesian Optimization module is introduced to better handle scenarios with noisy objective functions, particularly in image reconstruction-driven optimization tasks. The second optimization strategy relies on detailed GEANT4-based simulations, which, while more computationally intensive, offer higher physical fidelity. These simulations are also employed to study the impact of incorporating secondary particle information alongside cosmic muons for improved material discrimination. This paper highlights the outcomes and key findings from these optimization studies.
@@ I Introducton
Muon tomography, also known as muography, is a non-invasive imaging technique that leverages naturally occurring cosmic-ray muons to probe the internal structure of large objects. These muons originate from interactions between high-energy cosmic rays and atomic nuclei in the upper atmosphere, producing showers of secondary particles, primarily pions and kaons. As these mesons decay, they generate muons that travel toward the Earth’s surface with minimal energy loss. At sea level, the flux of cosmic-ray muons is approximately , with an average energy of vertical muons of the order . Due to their relatively large mass—roughly times that of an electron - these highly energetic muons have a strong penetrating ability, allowing them to traverse several meters of dense material. This makes them particularly valuable for imaging applications where conventional radiation-based methods, such as X-ray imaging, are ineffective due to limited penetration depth [Borozdin and others, 2003].
@@
Muon tomography can be classified into two main approaches: muon transmission imaging and muon scattering tomography (MST). The transmission method relies on detecting attenuation in muon flux as particles pass through an object, providing insights into its internal composition. This technique is particularly useful for large-scale applications, such as studying volcanic structures [D’Alessandro et al., 2018] or uncovering hidden chambers in archaeological sites [Alvarez and others, 1970]. However, it typically requires prolonged data acquisition periods, which makes it less suitable for applications where rapid imaging is essential. On the other hand, muon scattering tomography (MST) operates by measuring the deflections of muon trajectories caused by multiple Coulomb scattering interactions with the material they pass through. High-atomic-number (Z) materials, such as lead and uranium, tend to cause more significant scattering than lower-Z materials like aluminum or plastic. This property makes MST particularly useful for detecting dense and high-Z substances in diverse applications that span many fields, including nuclear reactor inspection, civil engineering, archaeology, and various industrial and homeland security uses [International Atomic Energy Agency, 2022]. In border security applications, MST is a powerful tool for identifying hidden threats and dense contraband in cargo containers, such as radioactive materials or explosives. Furthermore, MST is well-suited for scanning smaller objects and offers the advantage of rapid data acquisition, making it an ideal technology for real-time security applications where both speed and precision are critical for border security operations [Barnes et al., 2023].
@@
As muons traverse a material medium, they primarily interact with material nuclei through multiple Coulomb scatterings, where small-angle deviations are most frequent, while occasional large-angle deflections occur as described by Rutherford scattering. After crossing a macroscopic distance, the overall scattering angle distribution consists of a Gaussian core, which accounts for the majority of deflections due to the central limit theorem, along with a longer tail from less frequent large-angle scatterings. For a muon with momentum [MeV] passing through a material of radiation length [m] over a distance [m], the root mean square (RMS) width of the Gaussian component of the scattering angle distribution can be approximated by [Navas and others, 2024]:
@@
Optimizing the design parameters of an MST scanner is crucial to achieving optimal performance, as different configurations impact key factors such as acceptance, detection efficiency, and angular resolution. Design parameters such as the spatial arrangement and size of detectors play a critical role in determining the system’s ability to accurately track muon trajectories, maintain a high signal-to-noise ratio, and effectively classify materials. A suboptimal design may lead to reduced tracking precision, lower detection rates, and decreased overall scanning effectiveness. Prior optimization of MST systems has primarily relied on simulations combined with grid-search approaches to evaluate discrete configurations. However, a systematic, end-to-end optimization strategy that directly links detector design to material discrimination or scanning performance has, to our knowledge, not yet been explored. This motivates the approach presented in this work, which seeks to optimize the scanner design in a holistic, performance-driven manner.
In this work, we focus on optimizing the SilentBorder scanner (https://silentborder.eu/), a modular MST system designed for cargo screening applications. The scanner consists of multiple hodoscopes, each of which is a self-contained detector module comprising three equally-spaced sensitive detector panels inside a protective casing. The panels are made of a mat of plastic scintillator fibers 1 mm in diameter with a pitch of 7.5 mm, achieving a spatial resolution of 120 and an angular resolution of a few tenths of a milliradian at meter-scale separations [Barnes et al., 2023]. Compared to drift tubes scanners- for example, those used at LANL [Priedhorsky et al., 2003]- which achieve sub-millimeter resolutions but require complex infrastructure including gas handling, high-voltage systems, and bulky mechanical support, scintillating fibers offer a more compact solution at the cost of a high channel density and demanding optical/electronic readout. In this sense, fibers represent a balance between operational complexity and intrinsic spatial granularity. These hodoscopes are arranged around the scanned cargo, forming a tracking system capable of measuring muon trajectories before and after interaction with the inspected volume. Due to its modular design, the SilentBorder scanner offers flexibility in configuration, allowing different setups to be explored for improved detection performance. To systematically optimize the design parameterization of the hodoscopes, we adopt a dual-method approach combining Monte Carlo simulations and differentiable optimization techniques [Dorigo and others, 2023,Aehle and others, 2023]:
GEANT4-based simulation framework: We employ a GEANT4 [Agostinelli and others, 2003] simulation framework to model different SilentBorder scanner configurations and various cargo materials. GEANT4, a widely used toolkit for simulating particle interactions with matter, provides a realistic modeling environment for evaluating the performance of different scanner layouts.
TomOpt-based optimization: TomOpt [Strong and others, 2024] is a differentiable simulation framework developed for optimizing MST scanner designs through gradient descent.
@@ II GEANT4 Simulation Framework
A GEANT4-based framework was developed by the SilentBorder consortium to provide a simplified model of the SilentBorder scanner concept with the flexibility to simulate a variety of different scanner configurations and cargo contents.
Geometry Description Markup Language (GDML) [Chytracek et al., 2006] is used to describe the size, shape, and location of the relevant objects and materials. Firstly, an empty shipping container of external dimensions 6.07 m2.44 m2.60 m is placed at the center of the geometry with a plywood floor of depth 40 cm. The container is surrounded by a total of 48 hodoscopes, a 34 grid on each of the long sides of the container. This arrangement is visualized in Fig. 1. Each hodoscope is itself comprised of 3 layers (plates) of solid plastic scintillator each of dimensions 2.1 m1.1 m3.3 mm and separated by 10 cm. Note the casing and electronics for each hodoscope are not simulated in this simplification, but the 10 cm gaps are introduced to account for the ‘dead regions’ they introduce. The plate is simulated as a solid scintillator rather than as a mat of scintillator fibers. This significantly reduces the complexity of the geometry description and can be considered a good approximation of the fiber mats once the smearing for detector resolution effects is applied later.
@@
Air shower cosmic-ray muons can be simulated in this volume using the Muons for Silent Border (MuSiBo) generator. MuSiBo utilizes a Gaisser parameterization of the muon flux, with modifications to account for muon decay and the Earth’s curvature. It is based on work presented in reference [Blackwell and Kudryavtsev, 2015]. Muons are sampled on the surface of a cube, a method which better accounts for the high-angle flux when compared to sampling on a single surface above the region of interest. For this simulation, a box of 11 m5 m4 m was used to fully enclose the volume of the container and all of the hodoscopes.
Finally, the initial muons simulated with MuSiBo are propagated through the simulated geometry via use of the GEANT4 software package [Agostinelli and others, 2003]. Specifically, the QGSP_BERT physics list is utilized in version 11.0.p03 of the software. Muons, and any secondary particles produced, are stepped through the full geometry with information saved about their trajectory and total energy deposition within each logical volume.
@@ III GEANT4 Study of Hodoscope Positioning
The relative positioning of hodoscope plates can impact the performance of the scanner system. This impact can be appreciated in both the efficiency in collecting ‘interesting’ muons and the angular resolution of their trajectories.
A study is performed using the GEANT4 simulation framework described in Section II. For the rest of this section we will use a coordinate system that is relative to each hodoscope group; will refer to the long in-plane dimension, to the short in-plane dimension, and to the out-of-plane dimension. This ensures that moving hodoscope plates in one coordinate is conceptually the same regardless of the rotational orientation of the hodoscope group (top, bottom or sides).
@@
A baseline configuration was established with the container filled entirely with wood and the hodoscopes as close together in and as possible. Note that due to space reserved for electronics and casing this still leaves a 10 cm gap between the active scintillator areas of adjacent hodoscopes. The default spacing in between the three plates that form a single hodoscope is 10 cm.
@@
A number of scenarios are created by independently varying the spacing in , and . Due to physical constraints (such as not overlapping any hodoscope regions) the variations are different for each direction. The cargo is kept identical and for each scenario a 15 minute exposure to the cosmic flux is provided by the MuSiBo generator. This 15 minute exposure corresponds to a total of 12,683,700 muons in the volume used by MuSiBo, and the GEANT4 simulation required an average of 14.5 CPU hours to propagate and save information for each of these muons.
The energy depositions recorded by GEANT4 in the active hodoscope plates pass through a pseudo detector simulation. At this stage, in order to reduce computational time, we only consider events with 2 or more depositions before and 2 or more depositions after the cargo. Without this requirement a muon will not be reconstructed later on, so we can save resources by discarding this event here. This step is later referred to as ‘filtering’. Any depositions recorded within 1 mm and 20 ns of each other are combined and a Gaussian positional smearing, with width 1 mm, is then applied to their computed barycenter. This smearing, larger than the real intrinsic per-plate resolution of 120 m, is applied to conservatively approximate the combined effects of clustering, optical cross-talk, and readout not explicitly modeled in GEANT4. Finally, a threshold of 0.4 MeV is applied to ensure the deposition would have created a detector ‘hit’.
These hits form the input to a simple reconstruction chain. Tracks are formed when one hit is found on at least two hodoscope plates. These two plates must be found in different layers of the same hodoscope group (above, below, left or right). If a hodoscope plate has received multiple hits it is excluded from track building. This only impacts about 3.5% of track candidates, and therefore the small reduction in efficiency is worth accepting for the improved angular resolution. Whilst a more sophisticated algorithm could be designed to make use of this subset, the reduction in operating times for field-based MST scanners would be minimal.
@@
Linear regression is used to find the best fit line between the hits providing the direction of the reconstructed muon track. The final stage of the reconstruction takes events in which the muon had tracks reconstructed in two different hodoscope groups. A simple Point-of-Closest-Approach (PoCA) reconstruction is performed from which the location and angle of the scatter point is determined. An example of the output of the PoCA reconstruction is shown in Fig. 2. The reconstruction chain is significantly less computationally expensive than the simulation, each scenario requiring an average of 30 CPU minutes to perform the full reconstruction and save the output.
@@
The number of events in which a PoCA point was reconstructed.
The number of events in which a PoCA point was reconstructed inside the container - thus providing information about the cargo.
@@
In both directions the number of muons of interest increase as the gaps increase, this is purely because the hodoscopes as an ensemble now cover a larger volume within the total simulation world and hence more of the muons are likely to intercept at least one hodoscope. However, the number of PoCA points, and specifically those in the cargo, decreases due to the gaps that are created in the regions closest to the cargo itself. There is a subsequent reduction in the number of muons that intercept hodoscopes before and after the container and pass through the cargo. The reduction in useful PoCA points per muon will mean that a longer exposure time is required in order to achieve the same level of sensitivity.
@@
There is no visible change in the standard deviation of the angular difference metric, so the angular resolution does not improve. There is a small increase in the mean when going to the largest gaps. This is likely due to the reduced available phase space in the scattering angle created by the gaps.
@@
Fig. 5 shows the variation in efficiency and angular resolution metrics as a function of the vertical spacing between the hodoscope plates in . The nominal gap between the plates is 10 cm. As with the variations, the number of interesting muons increases as the total scanner volume increases with larger spacing. The number of PoCA points slightly decreases. This occurs because the probability of the muon hitting enough plates to ensure track reconstruction before and after the cargo decreases, but the effect is very minimal. The angular resolution on the other hand is dramatically impacted by the widening of the vertical spacing. By increasing the spacing both the standard deviation and mean parameters reduce significantly before beginning to flatten out. As the distance between the two or three detection points increases the positional resolution of each plate has a smaller impact on the reconstructed angle, hence improving the angular resolution. The reduction in the mean occurs due to the asymmetry of the distribution. As the resolution improves the asymmetry is naturally reduced in magnitude.
@@
These variation studies can be used to inform decisions about the design for a full-scale SilentBorder-like scanner system. As could be anticipated, minimizing the in-plane gaps between hodoscopes improves the efficiency of the system, thus reducing the necessary exposure time and making the system quicker to use and more viable in border security. Increasing the vertical out-of-plane spacing between hodoscope plates improves the angular resolution due to the lever arm effect, although it does so at the expense of efficiency. For a particular system an appropriate trade-off balance can be found, in the scenario examined in this study Fig. 5 suggests that up to 20 cm the increase in out-of-plane spacing improves the angular resolution enough for the reduction in efficiency to be worthwhile. Beyond this, the resolution improvements plateau and therefore would be unlikely to be worth the efficiency reduction. Note that these studies were completed with a simple PoCA reconstruction and decisions were made to ignore plates with multiple hits or events with 3 or more tracks. These considerations will have an impact on the magnitude of the effects presented but should not undermine the trajectories of the changes and thus the conclusions drawn.